2) A mass of 4 kg is moving at 9 m/s in the +x direction which collides inelastically with a mass (mass 2) of 7 kg. After the collision, the masses are moving at 3 m/s in the +x direction. What is the average force on mass 2 in Newtons if the collision takes 0.1 seconds? If in the negative x direction, include a negative sign.
Answer #1

here,

2)

m1 = 4 kg

m2 = 7 kg

initial velocity of 1 , u1 = 9 m/s

the final velocity , v = 3 m/s

let the initial velocity of 2 be u2

using conservation of momentum

m1 *u1 + m2 * u2 = (m1 + m2) * v

4 * 9 + 7 * u2 = ( 4 + 7) * 3

u2 = - 0.43 m/s

the initial velocity of 2 is - 0.43 m/s in -x direction

t = 0.1 s

the average force on m2 , F2 = m2 * ( v - u) /t

F2 = 7 * (3 - (-0.43)) /0.1 N

F2 = 520.1 N